One reference states that γ for trans-1,2-dichloroethene is 1.12. Calculate the percent error between your value and the reference value.
Answer #1

Given molecule is trans-1,2-dichloroethane as:

which is a non-linear molecule.

Total number of atoms = 6

So, total number of degrees f freedom = 6 *3 = 18

Number of translational degrees of freedom = 3

Number of rotational degrees of freedom = 3

So, the number of vibrational degrees of freedom = 18 - 3-3 = 12

Total energy (E) = 3 *1/2 RT + 3*1/2RT + 12 *RT = 15RT

Now, Cv = (E/ T)v = 15R

So, Cp = Cv + R = 15 R + R = 16 R

Hence, = Cp / Cv = 16R / 15 R = 1.07

But Value From reference = 1.12

Percent Error = | Calculated value - Reference Value| *100% / Reference value = (| 1.07 - 1.12| *100%) / 1.12 = 4.46 %